OpenStudy (meilendurcer):

5/8 - 3/10= a. 13/40 b. 1/20 c. 1/4 d. 1/5 i tink it's D

OpenStudy (anonymous):

\[\frac{ 5 }{ 8 } = \frac{ 50 }{ 80 }\] \[\frac{ 3 }{ 10 } = \frac{ 24 }{ 80 }\] \[50-24=26\] \[\frac{ 26 }{ 80 } = \frac{ 13 }{ 40 }\]

OpenStudy (meilendurcer):

@hkim how did you get that?

OpenStudy (anonymous):

Which part don't you get?

OpenStudy (meilendurcer):

the first two. i am trying another one like it so thats why.

OpenStudy (anonymous):

I just found the common denominator. The first denominator is 8, and the second denominator is 10. If you just multiply by the denominator of the other, so if you have a 5/8 and multiply it by 10, then multiply 3/10 and multiply it by 8, then you should get the denominator the same for both.